# SENTRA: Selected-Next-Token Transformer for LLM Text Detection

### v1

@@ 4.1 Datasets
If we define text similar to the training data distribution as in-domain and text that is dissimilar as out-of-domain, it is well established supervised LLM detection methods perform significantly better in-domain than out-of-domain [Dugan et al.(2024)Dugan, Hwang, Trhlík, Zhu, Ludan, Xu, Ippolito, and Callison-Burch]. However, a model designed for LLM-generated text detection in real world scenarios will inevitably encounter out-of-domain texts. For this reason, this work focuses on out-of-domain experiments, where key subsets of data are withheld from the training dataset.
To evaluate the effectiveness of our proposed method, we used three publicly available datasets: RAID [Dugan et al.(2024)Dugan, Hwang, Trhlík, Zhu, Ludan, Xu, Ippolito, and Callison-Burch], M4GT [Wang et al.(2024a)Wang, Mansurov, Ivanov, Su, Shelmanov, Tsvigun, Mohammed Afzal, Mahmoud, Puccetti, Arnold, Aji, Habash, Gurevych, and Nakov] and MAGE [Li et al.(2024a)Li, Li, Cui, Bi, Wang, Wang, Yang, Shi, and Zhang], focusing exclusively on English-language data.
@@
M4GT: An extension of M4 [Wang et al.(2024b)Wang, Mansurov, Ivanov, Su, Shelmanov, Tsvigun, Whitehouse, Mohammed Afzal, Mahmoud, Sasaki, Arnold, Aji, Habash, Gurevych, and Nakov], the M4GT dataset is a multi-domain and multi-LLM-generator corpus comprising data from domains, LLMs, and different detection tasks.
MAGE: The MAGE dataset covers content domains, with data generated by LLMs using different prompting strategies. It is specifically designed to assess out-of-distribution generalization capability. We use the "Unseen Domains" evaluation from [Li et al.(2024a)Li, Li, Cui, Bi, Wang, Wang, Yang, Shi, and Zhang].
Each dataset is further split into training, validation and test sets. For MAGE, we used the published split. To mitigate the label imbalance problem, the train and validation splits are balance-sampled to ensure an equal number of human- and LLM-generated texts. This was achieved by down-sampling the majority class to match the size of the minority class within split. Addressing this imbalance is crucial for two reasons: 1) the percentage of LLM-generated text is over in the RAID dataset by design; 2) across the three datasets, the proportion of LLM-generated text varies significantly. The average train and validation set sizes show how much data went into the training of the supervised methods while ensuring class balance, providing a clear comparison to the total dataset size. The MAGE dataset has significantly shorter texts and this adds difficulties to the detection task [Tian et al.(2024)Tian, Chen, Wang, Bai, ZHANG, Li, Xu, and Wang, Fraser et al.(2024)Fraser, Dawkins, and Kiritchenko].
@@ 4.2 Baseline Methods
We evaluated and compared the performance of our approach against multiple existing methods, including zero-shot, embedding-based, and supervised detectors. For zero-shot, we selected perplexity [Guo et al.(2023)Guo, Zhang, Wang, Jiang, Nie, Ding, Yue, and Wu], Fast-DetectGPT [Bao et al.(2024)Bao, Zhao, Teng, Yang, and Zhang], and Binoculars [Hans et al.(2024)Hans, Schwarzschild, Cherepanova, Kazemi, Saha, Goldblum, Geiping, and Goldstein] detectors. For embedding-based detectors, we selected UAR [Soto et al.(2024)Soto, Koch, Khan, Chen, Bishop, and Andrews] and evaluated both its Multi-LLM and Multi-domain models. For supervised detectors, we chose RoBERTa [Liu et al.(2019)Liu, Ott, Goyal, Du, Joshi, Chen, Levy, Lewis, Zettlemoyer, and Stoyanov] with direct fine-tuning, Ghostbuster [Verma et al.(2024)Verma, Fleisig, Tomlin, and Klein] which trains a logistic regression classifier on forward-selected crafted log-probability features, and Text Fluoroscopy [Yu et al.(2024)Yu, Chen, Yang, Zhang, and Yu] which utilizes intrinsic features. For RoBERTa, we used the same settings as the fine-tuning of SENTRA: a learning rate of , a weight decay of , and a patience of two epochs.
@@
We compared the baseline methods mentioned above with our proposed methods. We present results from two SENTRA encoder variations, R-SENTRA and SENTRA. R-SENTRA has all non-LLM weights in SENTRA encoder initialized at random (without pre-training), whereas the full SENTRA model has those weights pre-trained
as described in Section 3.3.
@@
Interestingly, prompting an LLM to do the LLM-text detection task is not well studied and does not appear in standard benchmarking work [Dugan et al.(2024)Dugan, Hwang, Trhlík, Zhu, Ludan, Xu, Ippolito, and Callison-Burch, Wang et al.(2024b)Wang, Mansurov, Ivanov, Su, Shelmanov, Tsvigun, Whitehouse, Mohammed Afzal, Mahmoud, Sasaki, Arnold, Aji, Habash, Gurevych, and Nakov, Li et al.(2024a)Li, Li, Cui, Bi, Wang, Wang, Yang, Shi, and Zhang]. We performed a small case study to evaluate how a SOTA LLM, GPT-4o [OpenAI et al.(2024a)OpenAI, Hurst, Lerer, Goucher, Perelman, Ramesh, Clark, Ostrow, Welihinda, Hayes, Radford, Mądry, Baker-Whitcomb, Beutel, Borzunov, Carney, Chow, Kirillov, Nichol, Paino, Renzin, Passos, Kirillov, Christakis, Conneau, Kamali, Jabri, Moyer, Tam, Crookes, Tootoochian, Tootoonchian, Kumar, Vallone, Karpathy, Braunstein, Cann, Codispoti, Galu, Kondrich, Tulloch, Mishchenko, Baek, Jiang, Pelisse, Woodford, Gosalia, Dhar, Pantuliano, Nayak, Oliver, Zoph, Ghorbani, Leimberger, Rossen, Sokolowsky, Wang, Zweig, Hoover, Samic, McGrew, Spero, Giertler, Cheng, Lightcap, Walkin, Quinn, Guarraci, Hsu, Kellogg, Eastman, Lugaresi, Wainwright, Bassin, Hudson, Chu, Nelson, Li, Shern, Conger, Barette, Voss, Ding, Lu, Zhang, Beaumont, Hallacy, Koch, Gibson, Kim, Choi, McLeavey, Hesse, Fischer, Winter, Czarnecki, Jarvis, Wei, Koumouzelis, Sherburn, Kappler, Levin, Levy, Carr, Farhi, Mely, Robinson, Sasaki, Jin, Valladares, Tsipras, Li, Nguyen, Findlay, Oiwoh, Wong, Asdar, Proehl, Yang, Antonow, Kramer, Peterson, Sigler, Wallace, Brevdo, Mays, Khorasani, Such, Raso, Zhang, Lohmann, Sulit, Goh, Oden, Salmon, Starace, Brockman, Salman, Bao, Hu, Wong, Wang, Schmidt, Whitney, Jun, Kirchner, Pinto, Ren, Chang, Chung, Kivlichan, O’Connell, O’Connell, Osband, Silber, Sohl, Okuyucu, Lan, Kostrikov, Sutskever, Kanitscheider, Gulrajani, Coxon, Menick, Pachocki, Aung, Betker, Crooks, Lennon, Kiros, Leike, Park, Kwon, Phang, Teplitz, Wei, Wolfe, Chen, Harris, Varavva, Lee, Shieh, Lin, Yu, Weng, Tang, Yu, Jang, Candela, Beutler, Landers, Parish, Heidecke, Schulman, Lachman, McKay, Uesato, Ward, Kim, Huizinga, Sitkin, Kraaijeveld, Gross, Kaplan, Snyder, Achiam, Jiao, Lee, Zhuang, Harriman, Fricke, Hayashi, Singhal, Shi, Karthik, Wood, Rimbach, Hsu, Nguyen, Gu-Lemberg, Button, Liu, Howe, Muthukumar, Luther, Ahmad, Kai, Itow, Workman, Pathak, Chen, Jing, Guy, Fedus, Zhou, Mamitsuka, Weng, McCallum, Held, Ouyang, Feuvrier, Zhang, Kondraciuk, Kaiser, Hewitt, Metz, Doshi, Aflak, Simens, Boyd, Thompson, Dukhan, Chen, Gray, Hudnall, Zhang, Aljubeh, Litwin, Zeng, Johnson, Shetty, Gupta, Shah, Yatbaz, Yang, Zhong, Glaese, Chen, Janner, Lampe, Petrov, Wu, Wang, Fradin, Pokrass, Castro, Castro, Pavlov, Brundage, Wang, Khan, Murati, Bavarian, Lin, Yesildal, Soto, Gimelshein, Cone, Staudacher, Summers, LaFontaine, Chowdhury, Ryder, Stathas, Turley, Tezak, Felix, Kudige, Keskar, Deutsch, Bundick, Puckett, Nachum, Okelola, Boiko, Murk, Jaffe, Watkins, Godement, Campbell-Moore, Chao, McMillan, Belov, Su, Bak, Bakkum, Deng, Dolan, Hoeschele, Welinder, Tillet, Pronin, Tillet, Dhariwal, Yuan, Dias, Lim, Arora, Troll, Lin, Lopes, Puri, Miyara, Leike, Gaubert, Zamani, Wang, Donnelly, Honsby, Smith, Sahai, Ramchandani, Huet, Carmichael, Zellers, Chen, Chen, Nigmatullin, Cheu, Jain, Altman, Schoenholz, Toizer, Miserendino, Agarwal, Culver, Ethersmith, Gray, Grove, Metzger, Hermani, Jain, Zhao, Wu, Jomoto, Wu, Shuaiqi, Xia, Phene, Papay, Narayanan, Coffey, Lee, Hall, Balaji, Broda, Stramer, Xu, Gogineni, Christianson, Sanders, Patwardhan, Cunninghman, Degry, Dimson, Raoux, Shadwell, Zheng, Underwood, Markov, Sherbakov, Rubin, Stasi, Kaftan, Heywood, Peterson, Walters, Eloundou, Qi, Moeller, Monaco, Kuo, Fomenko, Chang, Zheng, Zhou, Manassra, Sheu, Zaremba, Patil, Qian, Kim, Cheng, Zhang, He, Zhang, Jin, Dai, and Malkov], and a reasoning model, o1 [OpenAI et al.(2024b)OpenAI, Jaech, Kalai, Lerer, Richardson, El-Kishky, Low, Helyar, Madry, Beutel, Carney, Iftimie, Karpenko, Passos, Neitz, Prokofiev, Wei, Tam, Bennett, Kumar, Saraiva, Vallone, Duberstein, Kondrich, Mishchenko, Applebaum, Jiang, Nair, Zoph, Ghorbani, Rossen, Sokolowsky, Barak, McGrew, Minaiev, Hao, Baker, Houghton, McKinzie, Eastman, Lugaresi, Bassin, Hudson, Li, Bourcy, Voss, Shen, Zhang, Koch, Orsinger, Hesse, Fischer, Chan, Roberts, Kappler, Levy, Selsam, Dohan, Farhi, Mely, Robinson, Tsipras, Li, Oprica, Freeman, Zhang, Wong, Proehl, Cheung, Mitchell, Wallace, Ritter, Mays, Wang, Such, Raso, Leoni, Tsimpourlas, Song, Lohmann, Sulit, Salmon, Parascandolo, Chabot, Zhao, Brockman, Leclerc, Salman, Bao, Sheng, Andrin, Bagherinezhad, Ren, Lightman, Chung, Kivlichan, O’Connell, Osband, Gilaberte, Akkaya, Kostrikov, Sutskever, Kofman, Pachocki, Lennon, Wei, Harb, Twore, Feng, Yu, Weng, Tang, Yu, Candela, Palermo, Parish, Heidecke, Hallman, Rizzo, Gordon, Uesato, Ward, Huizinga, Wang, Chen, Xiao, Singhal, Nguyen, Cobbe, Shi, Wood, Rimbach, Gu-Lemberg, Liu, Lu, Stone, Yu, Ahmad, Yang, Liu, Maksin, Ho, Fedus, Weng, Li, McCallum, Held, Kuhn, Kondraciuk, Kaiser, Metz, Boyd, Trebacz, Joglekar, Chen, Tintor, Meyer, Jones, Kaufer, Schwarzer, Shah, Yatbaz, Guan, Xu, Yan, Glaese, Chen, Lampe, Malek, Wang, Fradin, McClay, Pavlov, Wang, Wang, Murati, Bavarian, Rohaninejad, McAleese, Chowdhury, Chowdhury, Ryder, Tezak, Brown, Nachum, Boiko, Murk, Watkins, Chao, Ashbourne, Izmailov, Zhokhov, Dias, Arora, Lin, Lopes, Gaon, Miyara, Leike, Hwang, Garg, Brown, James, Shu, Cheu, Greene, Jain, Altman, Toizer, Toyer, Miserendino, Agarwal, Hernandez, Baker, McKinney, Yan, Zhao, Hu, Santurkar, Chaudhuri, Zhang, Fu, Papay, Lin, Balaji, Sanjeev, Sidor, Broda, Clark, Wang, Gordon, Sanders, Patwardhan, Sottiaux, Degry, Dimson, Zheng, Garipov, Stasi, Bansal, Creech, Peterson, Eloundou, Qi, Kosaraju, Monaco, Pong, Fomenko, Zheng, Zhou, McCabe, Zaremba, Dubois, Lu, Chen, Cha, Bai, He, Zhang, Wang, Shao, and Li], could perform on a sample of the OOD datasets. We were unable, due to the high cost of these APIs, to run the full evaluation datasets through these models and therefore chose to randomly sample from the full datasets and perform a fair comparison on the smaller test sets. The evaluation results for the GPT4-o and o1 LLMs and their comparison with SENTRA performance are reported in Section 4.5.
@@ 4.3 Ablation Study
Table 1 shows the effect of pre-training SENTRA on all datasets. r-SENTRA is the "raw" SENTRA showing the architecture’s performance without pre-training on the M4GT dataset. Across the four datasets, the average and worst-case performance over the domains was increased after pre-training. This shows the contrastive pre-training method presented in Figure 2 is an effective method for improving SENTRA as an encoder for the LLM text detection.
Table 2 presents an ablation study on SENTRA components. Rows 2 and 3 of Table 2 show the AUROC performance metric after removing each of the two LLMs used to create SENTRA’s SNTP input (see Figure 1). Rows 4 and 5 of the table show the results when the Falcon-7b models [Almazrouei et al.(2023)Almazrouei, Alobeidli, Alshamsi, Cappelli, Cojocaru, Debbah, Goffinet, Hesslow, Launay, Malartic, Mazzotta, Noune, Pannier, and Penedo] are replaced by different pairs of LLMs: Qwen-2.5-3b [Qwen et al.(2025)Qwen, Yang, Yang, Zhang, Hui, Zheng, Yu, Li, Liu, Huang, Wei, Lin, Yang, Tu, Zhang, Yang, Yang, Zhou, Lin, Dang, Lu, Bao, Yang, Yu, Li, Xue, Zhang, Zhu, Men, Lin, Li, Tang, Xia, Ren, Ren, Fan, Su, Zhang, Wan, Liu, Cui, Zhang, and Qiu] and Gemma3-1b [Team et al.(2025)Team, Kamath, Ferret, Pathak, Vieillard, Merhej, Perrin, Matejovicova, Ramé, Rivière, Rouillard, Mesnard, Cideron, Grill, Ramos, Yvinec, Casbon, Pot, Penchev, Liu, Visin, Kenealy, Beyer, Zhai, Tsitsulin, Busa-Fekete, Feng, Sachdeva, Coleman, Gao, Mustafa, Barr, Parisotto, Tian, Eyal, Cherry, Peter, Sinopalnikov, Bhupatiraju, Agarwal, Kazemi, Malkin, Kumar, Vilar, Brusilovsky, Luo, Steiner, Friesen, Sharma, Sharma, Gilady, Goedeckemeyer, Saade, Feng, Kolesnikov, Bendebury, Abdagic, Vadi, György, Pinto, Das, Bapna, Miech, Yang, Paterson, Shenoy, Chakrabarti, Piot, Wu, Shahriari, Petrini, Chen, Lan, Choquette-Choo, Carey, Brick, Deutsch, Eisenbud, Cattle, Cheng, Paparas, Sreepathihalli, Reid, Tran, Zelle, Noland, Huizenga, Kharitonov, Liu, Amirkhanyan, Cameron, Hashemi, Klimczak-Plucińska, Singh, Mehta, Lehri, Hazimeh, Ballantyne, Szpektor, Nardini, Pouget-Abadie, Chan, Stanton, Wieting, Lai, Orbay, Fernandez, Newlan, Ji, Singh, Black, Yu, Hui, Vodrahalli, Greff, Qiu, Valentine, Coelho, Ritter, Hoffman, Watson, Chaturvedi, Moynihan, Ma, Babar, Noy, Byrd, Roy, Momchev, Chauhan, Sachdeva, Bunyan, Botarda, Caron, Rubenstein, Culliton, Schmid, Sessa, Xu, Stanczyk, Tafti, Shivanna, Wu, Pan, Rokni, Willoughby, Vallu, Mullins, Jerome, Smoot, Girgin, Iqbal, Reddy, Sheth, Põder, Bhatnagar, Panyam, Eiger, Zhang, Liu, Yacovone, Liechty, Kalra, Evci, Misra, Roseberry, Feinberg, Kolesnikov, Han, Kwon, Chen, Chow, Zhu, Wei, Egyed, Cotruta, Giang, Kirk, Rao, Black, Babar, Lo, Moreira, Martins, Sanseviero, Gonzalez, Gleicher, Warkentin, Mirrokni, Senter, Collins, Barral, Ghahramani, Hadsell, Matias, Sculley, Petrov, Fiedel, Shazeer, Vinyals, Dean, Hassabis, Kavukcuoglu, Farabet, Buchatskaya, Alayrac, Anil, Dmitry, Lepikhin, Borgeaud, Bachem, Joulin, Andreev, Hardin, Dadashi, and Hussenot]. From the results, we can see that Gemma3-1b [Team et al.(2025)Team, Kamath, Ferret, Pathak, Vieillard, Merhej, Perrin, Matejovicova, Ramé, Rivière, Rouillard, Mesnard, Cideron, Grill, Ramos, Yvinec, Casbon, Pot, Penchev, Liu, Visin, Kenealy, Beyer, Zhai, Tsitsulin, Busa-Fekete, Feng, Sachdeva, Coleman, Gao, Mustafa, Barr, Parisotto, Tian, Eyal, Cherry, Peter, Sinopalnikov, Bhupatiraju, Agarwal, Kazemi, Malkin, Kumar, Vilar, Brusilovsky, Luo, Steiner, Friesen, Sharma, Sharma, Gilady, Goedeckemeyer, Saade, Feng, Kolesnikov, Bendebury, Abdagic, Vadi, György, Pinto, Das, Bapna, Miech, Yang, Paterson, Shenoy, Chakrabarti, Piot, Wu, Shahriari, Petrini, Chen, Lan, Choquette-Choo, Carey, Brick, Deutsch, Eisenbud, Cattle, Cheng, Paparas, Sreepathihalli, Reid, Tran, Zelle, Noland, Huizenga, Kharitonov, Liu, Amirkhanyan, Cameron, Hashemi, Klimczak-Plucińska, Singh, Mehta, Lehri, Hazimeh, Ballantyne, Szpektor, Nardini, Pouget-Abadie, Chan, Stanton, Wieting, Lai, Orbay, Fernandez, Newlan, Ji, Singh, Black, Yu, Hui, Vodrahalli, Greff, Qiu, Valentine, Coelho, Ritter, Hoffman, Watson, Chaturvedi, Moynihan, Ma, Babar, Noy, Byrd, Roy, Momchev, Chauhan, Sachdeva, Bunyan, Botarda, Caron, Rubenstein, Culliton, Schmid, Sessa, Xu, Stanczyk, Tafti, Shivanna, Wu, Pan, Rokni, Willoughby, Vallu, Mullins, Jerome, Smoot, Girgin, Iqbal, Reddy, Sheth, Põder, Bhatnagar, Panyam, Eiger, Zhang, Liu, Yacovone, Liechty, Kalra, Evci, Misra, Roseberry, Feinberg, Kolesnikov, Han, Kwon, Chen, Chow, Zhu, Wei, Egyed, Cotruta, Giang, Kirk, Rao, Black, Babar, Lo, Moreira, Martins, Sanseviero, Gonzalez, Gleicher, Warkentin, Mirrokni, Senter, Collins, Barral, Ghahramani, Hadsell, Matias, Sculley, Petrov, Fiedel, Shazeer, Vinyals, Dean, Hassabis, Kavukcuoglu, Farabet, Buchatskaya, Alayrac, Anil, Dmitry, Lepikhin, Borgeaud, Bachem, Joulin, Andreev, Hardin, Dadashi, and Hussenot] is competitive with Falcon-7b, and could be an alternative for more compute constrained environments. These choices in LLMs are by no means an exhaustive search, and this ablation shows SENTRA can work with other LLM pairs while echoing Binocular’s result showing Falcon-7b is particularly effective [Hans et al.(2024)Hans, Schwarzschild, Cherepanova, Kazemi, Saha, Goldblum, Geiping, and Goldstein].

### v2

@@ 4.1 Datasets
If we define text similar to the training data distribution as in-domain and text that is dissimilar as out-of-domain, it is well established supervised LLM detection methods perform significantly better in-domain than out-of-domain Dugan et al. (2024). However, a model designed for LLM-generated text detection in real world scenarios will inevitably encounter out-of-domain texts. For this reason, this work focuses on out-of-domain experiments, where key subsets of data are withheld from the training dataset.
To evaluate the effectiveness of our proposed method, we used three publicly available datasets: RAID Dugan et al. (2024), M4GT Wang et al. (2024a) and MAGE Li et al. (2024a), focusing exclusively on English-language data.
@@
M4GT: An extension of M4 Wang et al. (2024b), the M4GT dataset is a multi-domain and multi-LLM-generator corpus comprising data from domains, LLMs, and different detection tasks.
MAGE: The MAGE dataset covers content domains, with data generated by LLMs using different prompting strategies. It is specifically designed to assess out-of-distribution generalization capability. We use the "Unseen Domains" evaluation from Li et al. (2024a).
Each dataset is further split into training, validation and test sets. For MAGE, we used the published split. To mitigate the label imbalance problem, the train and validation splits are balance-sampled to ensure an equal number of human- and LLM-generated texts. This was achieved by down-sampling the majority class to match the size of the minority class within split. Addressing this imbalance is crucial for two reasons: 1) the percentage of LLM-generated text is over in the RAID dataset by design; 2) across the three datasets, the proportion of LLM-generated text varies significantly. The average train and validation set sizes show how much data went into the training of the supervised methods while ensuring class balance, providing a clear comparison to the total dataset size. The MAGE dataset has significantly shorter texts and this adds difficulties to the detection task Tian et al. (2024); Fraser et al. (2024).
@@ 4.2 Baseline Methods
We evaluated and compared the performance of our approach against multiple existing methods, including zero-shot, embedding-based, and supervised detectors. For zero-shot, we selected perplexity Guo et al. (2023), Fast-DetectGPT Bao et al. (2024), and Binoculars Hans et al. (2024) detectors. For embedding-based detectors, we selected UAR Soto et al. (2024) and evaluated both its Multi-LLM and Multi-domain models. For supervised detectors, we chose RoBERTa Liu et al. (2019) with direct fine-tuning, Ghostbuster Verma et al. (2024) which trains a logistic regression classifier on forward-selected crafted log-probability features, and Text Fluoroscopy Yu et al. (2024) which utilizes intrinsic features. For RoBERTa, we used the same settings as the fine-tuning of SENTRA: a learning rate of , a weight decay of , and a patience of two epochs.
@@
We compared the baseline methods mentioned above with our proposed methods. We present results from two SENTRA encoder variations, R-SENTRA and SENTRA. R-SENTRA has all non-LLM weights in SENTRA encoder initialized at random (without pre-training), whereas the full SENTRA model has those weights pre-trained as described in Section 3.3.
@@
Interestingly, prompting an LLM to do the LLM-text detection task is not well studied and does not appear in standard benchmarking work Dugan et al. (2024); Wang et al. (2024b); Li et al. (2024a). We performed a small case study to evaluate how a SOTA LLM, GPT-4o OpenAI et al. (2024a), and a reasoning model, o1 OpenAI et al. (2024b), could perform on a sample of the OOD datasets. We were unable, due to the high cost of these APIs, to run the full evaluation datasets through these models and therefore chose to randomly sample from the full datasets and perform a fair comparison on the smaller test sets. The evaluation results for the GPT4-o and o1 LLMs and their comparison with SENTRA performance are reported in Section 4.5.
@@ 4.3 Ablation Study
Table 1 shows the effect of pre-training SENTRA on all datasets. r-SENTRA is the "raw" SENTRA showing the architecture’s performance without pre-training on any dataset and then evaluating on the M4GT dataset. Across the four datasets, the average and worst-case performance over the domains was increased after pre-training. This shows the contrastive pre-training method presented in Figure 2 is an effective method for improving SENTRA as an encoder for the LLM text detection.
Table 2 presents an ablation study on SENTRA components. Rows 2 and 3 of Table 2 show the AUROC performance metric after removing each of the two LLMs used to create SENTRA’s SNTP input (see Figure 1). Rows 4 and 5 of the table show the results when the Falcon-7b models Almazrouei et al. (2023) are replaced by different pairs of LLMs: Qwen-2.5-3b Qwen et al. (2025) and Gemma3-1b Team et al. (2025). From the results, we can see that Gemma3-1b Team et al. (2025) is competitive with Falcon-7b, and could be an alternative for more compute constrained environments. These choices in LLMs are by no means an exhaustive search, and this ablation shows SENTRA can work with other LLM pairs while echoing Binocular’s result showing Falcon-7b is particularly effective Hans et al. (2024).